OpenStudy (anonymous):

On Friday, Linda performed 17 haircuts. This was 7 more than twice the number of haircuts, h, she performed on Monday. How many haircuts did Linda perform on Monday?

OpenStudy (anonymous):

You can represent this statement as the mathematical expression 17 = 7 + 2h and solve for h to get that she performed 5 haircuts on Monday.

OpenStudy (anonymous):

5*2=10+7

OpenStudy (anonymous):

take 17 subtract 7 =10 divide by 2 =5 5 haircuts monday